# Exact Diagonalization, Matrix Product States and Conformal Perturbation Theory Study of a 3D Ising Fuzzy Sphere Model

### v1

@@ 5.5 $V_{0}$ dependence of CPT couplings
The initial fuzzy sphere work for the 3D Ising CFT used a specific value of $V_{0}=4.75$ for most of their simulations Zhu2023 . In this section we discuss in what sense this particular choice of $V_{0}$ is fine-tuned to minimize finite size effects. We do this by analyzing ED data for two additional values of $V_{0}$, namely $V_{0}=2.5$ and $V_{0}=6$.
@@
Based on the analysis of the structure of the effects of perturbing operators on energy levels in the $\sigma$ and $\varepsilon$ families presented in Sec. 5.3, we devise the following scheme: We track the energy levels corresponding to $\sigma$, $\varepsilon$, $\partial\sigma$ and (=$\partial^{2}_{(\ell=2)}\sigma$, the idea being that those levels are affected almost exclusively by $g_{\varepsilon}$, $g_{\varepsilon^{\prime}}$ and $g_{C}$. [^19] We neglect here a smaller effect of the $[\sigma\sigma]_{0}$ family spin-6 operator on (=$\partial^{2}_{(\ell=2)}\sigma$. Using these four energy gaps we can therefore determine $v/R$, $g_{\varepsilon}$, $g_{\varepsilon^{\prime}}$ and $g_{C}$ using the corresponding first order CPT corrections to the CFT energy levels. Note that this is not a fit, as we have as many parameters as equations.
@@
Figure 11: By tracking the energy levels $\sigma$, $\varepsilon$, $\partial\sigma$ and (=$\partial^{2}_{(\ell=2)}\sigma$, we extract the couplings coefficients $g_{\varepsilon}$ (first column), $g_{\varepsilon^{\prime}}$ (second column) and $g_{C}$ (third column), for several values of $V_{0}$ (first row: $2.5$, second row $4.75$ and third row $6.0$) and varying transverse field $h$. The effective critical field corresponds to $g_{\varepsilon}=0$. For $V_{0}=4.75$ selected in Ref. (Zhu2023, ), the remaining $g_{\varepsilon^{\prime}}$ is one order of magnitude lower than for the other two $V_{0}$s, explaining the small observed finite size effects. $g_{C}$ is also significantly smaller, though presents in all cases peculiar, non-decreasing finite-size effects. The bottom line shows the scaling of the different couplings $g_{\cal V}$ off-criticality ($h=3$) as a function of the radius $R$ for the three $V_{0}$, compared to the expected behavior obtained from the CFT $R^{3-\Delta_{\cal V}}$. While in several cases, we observe a qualitative agreement, we observe several anomalies, e.g. $g_{\varepsilon}\propto R^{3-\Delta_{\varepsilon^{\prime}}}$ for $V_{0}=2.5$, that we discuss in Sec. 5.5.1.
@@
In Fig. 11 we perform this analysis for all three values of $V_{0}$ and plot the obtained coupling constants in a certain $h$ window around the critical point, indicated by the zero crossing of $g_{\varepsilon}$ as a function of $h$. We are then interested in analyzing the residual couplings when $g_{\varepsilon}=0$. For $V_{0}=2.5$ we see in the top row of Fig. 11 that the new coupling $g_{\varepsilon^{\prime}}$ is quite large and positive, with a value of $4\pi\times g_{\varepsilon^{\prime}}\approx 0.07$ for $N=16$. Furthermore, $g_{C}$ is also nonzero, but with a negative value instead. We can also see a monotonic decrease of the magnitude of these two subleading coupling constants as we increase the system size $N$, as expected from the renormalization group perspective. For the value $V_{0}=6$ we also observe sizable coupling constants in the lowest row of Fig. 11, however the signs of the coupling constants are now flipped compared to the $V_{0}=2.5$ case. The finite size behavior for $g_{C}$ is peculiar, as the extracted values seem to increase and then stagnate for the system sizes $N$ considered here. Finally in the middle row we show the application of this extended CPT model to the case $V_{0}=4.75$ considered before. The values of $g_{\varepsilon^{\prime}}$ are now an order of magnitude smaller than for the other two considered values of $V_{0}$. This finding highlights that the chosen value $V_{0}=4.75$ of Ref. Zhu2023 seems to correspond to fine tuning $g_{\varepsilon^{\prime}}\approx 0$, as conjectured in Lao:2023zis . (The same conclusion was also reached in Fardelli:2024qla .) The remnant coupling $g_{C}$ at the critical point is roughly a factor two smaller in magnitude compared to the other two values of $V_{0}$, but is not behaving monotonously with increasing system size (the same observation also applies for the residual small value of $g_{\varepsilon^{\prime}}$). So we see that despite the approximate vanishing of $g_{\varepsilon^{\prime}}$, there are still some residual couplings or unaccounted effects at work, which deserve a better understanding. We will next discuss some possible approaches.
@@ 5.5.1 Comments on the $N$ dependence of CPT couplings
The first potential reason is that there are higher-order effects in the RG running. For example, at the second order the beta-function equations would have the form (Cardy:1996xt, , p. 89)
@@
Could the second order term in the beta function explain the observed drift of $h_{c}$? The answer appears to be negative. To explain the drift, one needs an extra contribution to $4\pi g_{\varepsilon}$, compared to the overall rescaling by $(R_{2}/R_{1})^{2-\Delta_{\varepsilon}}$, where $R_{2}=\sqrt{16}$ and $R_{1}=\sqrt{6}$. This extra needed contribution is about $-0.2$ for $V_{0}=2.5$ and about $0.1$ for $V_{0}=6$. The extra contribution from the quadratic term is an order of magnitude smaller than this. Moreover it has the same sign for $V_{0}=2.5$ and $V_{0}=6$, since all couplings flip sign.
@@
The latter scenario can be in principle tested by going either to 1+1D or to the flat torus in 2+1D Whitsitt:2017ocl . In both cases curvature is zero and the drift should be absent (up to higher order RG effects). Indeed, we have not observed a drift of $h_{c}$ in the 1+1D case studied in Sec. 4. We leave 2+1D torus tests to the future.
@@ 6 Extraction of OPE coefficients $f_{\mathcal{O}\mathcal{O}\varepsilon}$
$f^{\rm ED}_{\Phi\Phi\varepsilon}\approx\begin{cases}f_{\mathcal{O}\mathcal{O}\varepsilon}&\Phi={\cal O}\text{ scalar primary}\,,\\ A(k,\rho)f_{\mathcal{O}\mathcal{O}\varepsilon}&\Phi\text{ level-$k$ descendantof a scalar primary ${\cal O}$.}\end{cases}$ (40)
@@
This procedure delivers rather accurate and insightful data presented in Fig. 12. The plot is arranged into four columns, the left two being the $\sigma$ family (primary and descendants up to level four) and the $\varepsilon$ family (primary and descendants up to level three). The third columns contains the the stress energy tensor primary and descendants up to level 2. The rightmost column contains a selection of other low-lying primaries and some their descendants. Spin-4 descendants of $\varepsilon$ (level 2) and of the stress tensor (level 2) are also reported in this column for reasons that will become clear. Based on our prescription, the OPEs for $\sigma$ and $\partial\sigma$ are the ones we used to determine $g_{\varepsilon}$ in the first place, and hence identically agree with CFT predictions. All the other OPE estimates reported in this figure however are independent ED data. The short horizontal lines in the left of the corresponding panels are $f^{\rm shift}_{\Phi\Phi\varepsilon}$ from other sources, either CB or prior fuzzy sphere work (see below).
@@
In the $\varepsilon$ family we observe that the primary and the large spin descendants exhibit similar behavior as discussed for the $\sigma$ family. However we observe strikingly large finite size effects for the OPE coefficients of the levels labeled (=$\partial^{2}_{(\ell=0)}\varepsilon$ and (=$\partial^{3}_{(\ell=1)}\varepsilon$. Our interpretation of this finding is that there is level repulsion, i.e. an off-diagonal CPT matrix element from other perturbations beyond $g_{\varepsilon}$ which couples the two energetically nearby levels (=$\partial^{2}_{(\ell=0)}\varepsilon$ and $\varepsilon^{\prime}$ and also (=$\partial^{3}_{(\ell=1)}\varepsilon$ and $\partial^{1}\varepsilon^{\prime}$. There was first evidence for this effect already in an avoided level crossing visible in the energy spectrum itself as shown in Fig. 8, most clearly between (=$\partial^{3}_{(\ell=1)}\varepsilon$ and $\partial^{1}\varepsilon^{\prime}$, and to a lesser extent between (=$\partial^{2}_{(\ell=0)}\varepsilon$ and $\varepsilon^{\prime}$. See Sec. 7 below for a detailed discussion.
@@
The third panel displays the OPE coefficients for the stress energy tensor $T^{\mu\nu}$ and some of its descendants. The OPE coefficient for the primary seems to converge smoothly towards the high precision rigorous CB result (89). [^20] An alternative method from Ref. Hu:2023xak required a bolder extrapolation in $N$ and produced a value somewhat below (89), see App. B.4. Our data supports (89) rather than the result of Ref. Hu:2023xak . The descendants exhibit larger finite size effects but do not seem to signal a tension with the expected OPE coefficients for the descendants of the stress energy tensor.
@@
In the last panel we show the extracted OPE coefficients for a few more primaries, such as $\varepsilon^{\prime}$, $C$, $\sigma^{\prime}$, $\sigma^{\mu\nu}$ and $\sigma^{\eta\mu\nu}$. While some of the considered levels have quite smooth and small finite size effects and will be discussed below, we see that $\varepsilon^{\prime}$ and $\partial\varepsilon^{\prime}$ show dramatically large effects. These are the respective partners of the levels discussed in the $\varepsilon$ family, which are coupled in the avoided level crossing scenario. And indeed they have the opposite behavior starting off from small apparent OPE values growing and then converging to OPE values of $2\sim 2.5$, while the two concerned levels in the $\varepsilon$ families show the opposite trend. They start at seemingly large OPE coefficients for small $N$ and then drop rapidly and start to level off towards OPE coefficients around $0.6\sim 0.7$. A natural interpretation is that at very small $N$ the order of two pairs of ED energy levels is inverted compared to the CFT expectation. Then at intermediate $N$ the two energy levels and their wave functions mix because of a large off-diagonal matrix element, while at large $N$ the expected CFT ordering is established and the mixing starts to fade out. We also observe a similar level reordering and accompanying avoided level crossings for three spin-4 states $C$, (=$\partial^{2}_{(\ell=4)}T_{\mu\nu}$ and (=$\partial^{4}_{(\ell=4)}\varepsilon$, however in this case the mixing matrix elements between the levels are much smaller, making the avoided level crossings apparent only in a much smaller $N$ range.
@@
The other primaries not involved in a level repulsion scenario exhibit quite small finite size effects and allow us to report some new OPE coefficients $f^{\rm shift}_{\sigma^{\prime}\sigma^{\prime}\varepsilon}\approx 2.90$, $f^{\rm shift}_{\sigma^{\mu\nu}\sigma^{\mu\nu}\varepsilon}\approx 1.46$ and $f^{\rm shift}_{\sigma^{\kappa\mu\nu}\sigma^{\kappa\mu\nu}\varepsilon}\approx 1.38$. We are not attempting a finite size extrapolation, we just plot the bona fide results from our numerical procedure. One of these coefficients was previously reported in Ref. Hu:2023xak as $f^{\rm shift}_{\sigma^{\prime}\sigma^{\prime}\varepsilon}\approx 2.98(13)$, measured in a different OPE extraction scheme in the fuzzy sphere setup, which yields exactly the same quantity in the $N\to\infty$ limit. We currently believe our result to be more accurate, as it does not rely on a similarly daunting extrapolation in $N$.
@@
In conclusion, the OPE extraction method which we presented shows smaller finite size corrections than the previous method Ref. Hu:2023xak . It is also relatively straightforward to implement. The main limitation of the method is that it is tailored to extracting OPE coefficients of the form $f_{\Phi\Phi\epsilon}$. We could also try to perturb the 3D Ising CFT by the $\sigma$ operator, which at the level of the microscopic model corresponds to weakly breaking the $\mathbb{Z}_{2}$ symmetry. However, the energy levels $E_{\Phi}$ will be corrected to second order in the $\sigma$ coupling, and these corrections would involve OPE coefficients $f_{\Phi\Psi\sigma}$ for all $\Psi$’s having the opposite $\mathbb{Z}_{2}$ quantum number from $\Phi$. So it’s not clear if these scheme could lead to a method extracting the OPE coefficients involving the $\sigma$ operator.

### v2

@@ 5.5 $V_{0}$ dependence of CPT couplings
The initial fuzzy sphere work for the 3D Ising CFT used a specific value of $V_{0}=4.75$ for most of their simulations [1]. In this section we discuss in what sense this particular choice of $V_{0}$ is fine-tuned to minimize finite size effects. We do this by analyzing ED data for two additional values of $V_{0}$, namely $V_{0}=2.5$ and $V_{0}=6$.
@@
Based on the analysis of the structure of the effects of perturbing operators on energy levels in the $\sigma$ and $\varepsilon$ families presented in Sec. 5.3, we devise the following scheme: We track the energy levels corresponding to $\sigma$, $\varepsilon$, $\partial\sigma$ and $\partial^{2}_{(\ell=2)}\sigma$, the idea being that those levels are affected almost exclusively by $g_{\varepsilon}$, $g_{\varepsilon^{\prime}}$ and $g_{C}$. [^22] We neglect here a smaller effect of the $\sigma\sigma]_{0}$ family spin-6 operator on $\partial^{2}_{(\ell=2)}\sigma$. Using these four energy gaps we can therefore determine $v/R$, $g_{\varepsilon}$, $g_{\varepsilon^{\prime}}$ and $g_{C}$ using the corresponding first order CPT corrections to the CFT energy levels. Note that this is not a fit, as we have as many parameters as equations.
@@
Figure 11: By tracking the energy levels $\sigma$, $\varepsilon$, $\partial\sigma$ and $\partial^{2}_{(\ell=2)}\sigma$, we extract the couplings coefficients $g_{\varepsilon}$ (first column), $g_{\varepsilon^{\prime}}$ (second column) and $g_{C}$ (third column), for several values of $V_{0}$ (first row: $2.5$, second row $4.75$ and third row $6.0$) and varying transverse field $h$. The effective critical field corresponds to $g_{\varepsilon}=0$. For $V_{0}=4.75$ selected in Ref. [1], the remaining $g_{\varepsilon^{\prime}}$ is one order of magnitude lower than for the other two $V_{0}$s, explaining the small observed finite size effects. $g_{C}$ is also significantly smaller, though presents in all cases peculiar, non-decreasing finite-size effects. The bottom line shows the scaling of the different couplings $g_{\cal V}$ off-criticality ($h=3$) as a function of the radius $R$ for the three $V_{0}$, compared to the expected behavior obtained from the CFT $R^{3-\Delta_{\cal V}}$. While in several cases, we observe a qualitative agreement, we observe several anomalies, e.g. $g_{\varepsilon}\propto R^{3-\Delta_{\varepsilon^{\prime}}}$ for $V_{0}=2.5$, that we discuss in Sec. 5.5.1.
@@
In Fig. 11 we perform this analysis for all three values of $V_{0}$ and plot the obtained coupling constants in a certain $h$ window around the critical point, indicated by the zero crossing of $g_{\varepsilon}$ as a function of $h$. We are then interested in analyzing the residual couplings when $g_{\varepsilon}=0$. For $V_{0}=2.5$ we see in the top row of Fig. 11 that the new coupling $g_{\varepsilon^{\prime}}$ is quite large and positive, with a value of $4\pi\times g_{\varepsilon^{\prime}}\approx 0.07$ for $N=16$. Furthermore, $g_{C}$ is also nonzero, but with a negative value instead. We can also see a monotonic decrease of the magnitude of these two subleading coupling constants as we increase the system size $N$, as expected from the renormalization group perspective. For the value $V_{0}=6$ we also observe sizable coupling constants in the lowest row of Fig. 11, however the signs of the coupling constants are now flipped compared to the $V_{0}=2.5$ case. The finite size behavior for $g_{C}$ is peculiar, as the extracted values seem to increase and then stagnate for the system sizes $N$ considered here. Finally in the middle row we show the application of this extended CPT model to the case $V_{0}=4.75$ considered before. The values of $g_{\varepsilon^{\prime}}$ are now an order of magnitude smaller than for the other two considered values of $V_{0}$. This finding highlights that the chosen value $V_{0}=4.75$ of Ref. [1] seems to correspond to fine tuning $g_{\varepsilon^{\prime}}\approx 0$, as conjectured in [29]. (The same conclusion was also reached in [22].) The remnant coupling $g_{C}$ at the critical point is roughly a factor two smaller in magnitude compared to the other two values of $V_{0}$, but is not behaving monotonously with increasing system size (the same observation also applies for the residual small value of $g_{\varepsilon^{\prime}}$). So we see that despite the approximate vanishing of $g_{\varepsilon^{\prime}}$, there are still some residual couplings or unaccounted effects at work, which deserve a better understanding. We will next discuss some possible approaches.
@@ 5.5.1 Comments on the $N$ dependence of CPT couplings
The first potential reason is that there are higher-order effects in the RG running. For example, at the second order the beta-function equations would have the form [55, p. 89]
@@
Could the second order term in the beta function explain the observed drift of $h_{c}$? The answer appears to be negative. To explain the drift, one needs an extra contribution to $4\pi g_{\varepsilon}$, compared to the overall rescaling by $(R_{2}/R_{1})^{2-\Delta_{\varepsilon}}$, where $R_{2}=\sqrt{16}$ and $R_{1}=\sqrt{6}$. This extra needed contribution is about 0.2-0.2 for $V_{0}=2.5$ and about $0.1$ for $V_{0}=6$. The extra contribution from the quadratic term is an order of magnitude smaller than this. Moreover it has the same sign for $V_{0}=2.5$ and $V_{0}=6$, since all couplings flip sign.
@@
The latter scenario can be in principle tested by going either to 1+1D or to the flat torus in 2+1D [64]. In both cases curvature is zero and the drift should be absent (up to higher order RG effects). Indeed, we have not observed a drift of $h_{c}$ in the 1+1D case studied in Sec. 4. We leave 2+1D torus tests to the future.
@@ 6 Extraction of OPE coefficients $f_{\mathcal{O}\mathcal{O}\varepsilon}$
[…]
@@
To use this procedure, we need to know how $g_{\epsilon}$ varies near the critical point. We use here the minimal fit which determines $g_{\varepsilon}$ and the speed of light from the first-order CPT for $\sigma$ and $\partial\sigma$ energy levels. The OPE coefficient $f_{\sigma\sigma\epsilon}$ used in this fit is assumed known and fixed to its CB value. [^23] One could alternatively say that this procedure determined ratios of other OPE coefficients divided by $f_{\sigma\sigma\epsilon}$.
This procedure delivers rather accurate and insightful data presented in Fig. 12. The plot is arranged into four columns, the left two being the $\sigma$ family (primary and descendants up to level four) and the $\varepsilon$ family (primary and descendants up to level three). The third columns contains the the stress energy tensor primary and descendants up to level 2. The rightmost column contains a selection of other low-lying primaries and some their descendants. Spin-4 descendants of $\varepsilon$ (level 2) and of the stress tensor (level 2) are also reported in this column for reasons that will become clear. Based on our prescription, the OPEs for $\sigma$ and $\partial\sigma$ are the ones we used to determine $g_{\varepsilon}$ in the first place, and hence identically agree with CFT predictions. All the other OPE estimates reported in this figure however are independent ED data. The short horizontal lines in the left of the corresponding panels are $f^{\rm shift}_{\Phi\Phi\varepsilon}$ from other sources, either conformal bootstrap (CB) or prior fuzzy sphere work (see below).
@@
In the $\varepsilon$ family we observe that the primary and the large spin descendants exhibit similar behavior as discussed for the $\sigma$ family. However we observe strikingly large finite size effects for the OPE coefficients of the levels labeled $\partial^{2}_{(\ell=0)}\varepsilon$ and $\partial^{3}_{(\ell=1)}\varepsilon$. Our interpretation of this finding is that there is level repulsion, i.e. an off-diagonal CPT matrix element from other perturbations beyond $g_{\varepsilon}$ which couples the two energetically nearby levels $\partial^{2}_{(\ell=0)}\varepsilon$ and $\varepsilon^{\prime}$ and also $\partial^{3}_{(\ell=1)}\varepsilon$ and $\partial^{1}\varepsilon^{\prime}$. There was first evidence for this effect already in an avoided level crossing visible in the energy spectrum itself as shown in Fig. 8, most clearly between $\partial^{3}_{(\ell=1)}\varepsilon$ and $\partial^{1}\varepsilon^{\prime}$, and to a lesser extent between $\partial^{2}_{(\ell=0)}\varepsilon$ and $\varepsilon^{\prime}$. See Sec. 7 below for a detailed discussion.
@@
The third panel displays the OPE coefficients for the stress energy tensor $T^{\mu\nu}$ and some of its descendants. The OPE coefficient for the primary seems to converge smoothly towards the high precision rigorous CB result (89). [^24] An alternative method from Ref. [14] required a bolder extrapolation in $N$ and produced a value somewhat below (89), see App. B.4. Our data supports (89) rather than the result of Ref. [14]. The descendants exhibit larger finite size effects but do not seem to signal a tension with the expected OPE coefficients for the descendants of the stress energy tensor.
@@
In the last panel we show the extracted OPE coefficients for a few more primaries, such as $\varepsilon^{\prime}$, $C$, $\sigma^{\prime}$, $\sigma^{\mu\nu}$ and $\sigma^{\eta\mu\nu}$. While some of the considered levels have quite smooth and small finite size effects and will be discussed below, we see that $\varepsilon^{\prime}$ and $\partial\varepsilon^{\prime}$ show dramatically large effects. These are the respective partners of the levels discussed in the $\varepsilon$ family, which are coupled in the avoided level crossing scenario. And indeed they have the opposite behavior starting off from small apparent OPE values growing and then converging to OPE values of $2\sim 2.5$, while the two concerned levels in the $\varepsilon$ families show the opposite trend. They start at seemingly large OPE coefficients for small $N$ and then drop rapidly and start to level off towards OPE coefficients around $0.6\sim 0.7$. A natural interpretation is that at very small $N$ the order of two pairs of ED energy levels is inverted compared to the CFT expectation. Then at intermediate $N$ the two energy levels and their wave functions mix because of a large off-diagonal matrix element, while at large $N$ the expected CFT ordering is established and the mixing starts to fade out. We also observe a similar level reordering and accompanying avoided level crossings for three spin-4 states $C$, $\partial^{2}_{(\ell=4)}T_{\mu\nu}$ and $\partial^{4}_{(\ell=4)}\varepsilon$, however in this case the mixing matrix elements between the levels are much smaller, making the avoided level crossings apparent only in a much smaller $N$ range.
@@
The other primaries not involved in a level repulsion scenario exhibit quite small finite size effects and allow us to report some new OPE coefficients $f^{\rm shift}_{\sigma^{\prime}\sigma^{\prime}\varepsilon}\approx 2.90$, $f^{\rm shift}_{\sigma^{\mu\nu}\sigma^{\mu\nu}\varepsilon}\approx 1.46$ and $f^{\rm shift}_{\sigma^{\kappa\mu\nu}\sigma^{\kappa\mu\nu}\varepsilon}\approx 1.38$. We are not attempting a finite size extrapolation, we just plot the bona fide results from our numerical procedure. One of these coefficients was previously reported in Ref. [14] as $f^{\rm shift}_{\sigma^{\prime}\sigma^{\prime}\varepsilon}\approx 2.98(13)$, measured in a different OPE extraction scheme in the fuzzy sphere setup, which yields exactly the same quantity in the $N\to\infty$ limit. We currently believe our result to be more accurate, as it does not rely on a similarly daunting extrapolation in $N$.
@@
In conclusion, the OPE extraction method which we presented shows smaller finite size corrections than the previous method Ref. [14]. It is also relatively straightforward to implement. The main limitation of the method is that it is tailored to extracting OPE coefficients of the form $f_{\Phi\Phi\epsilon}$. We could also try to perturb the 3D Ising CFT by the $\sigma$ operator, which at the level of the microscopic model corresponds to weakly breaking the $\mathbb{Z}_{2}$ symmetry. However, the energy levels $E_{\Phi}$ will be corrected to second order in the $\sigma$ coupling, and these corrections would involve OPE coefficients $f_{\Phi\Psi\sigma}$ for all $\Psi$’s having the opposite $\mathbb{Z}_{2}$ quantum number from $\Phi$. So it’s not clear if these scheme could lead to a method extracting the OPE coefficients involving the $\sigma$ operator.